User Tools

Site Tools


handover:a1504

A1504

Hb HOB+0127 12001.380 GB

19:00 UT - Experiment auto-started OK (vk)

20:48 UT - Schedule halted for spaceX boresight check (vk)

22:06 UT - Schedule resumed. Missed scans: 039-2047 to 039-2205b. 192.5 GB deficit.

05:45 UT - Monica.Rxmon restarted (Liza)

12:07 UT - Alarm no PCFS activity for 15 min, due to there being no space for the log file (JS)

12:30 UT - Jamie “deleted some stuff” and this sorted the issue (JS)

Ke HOB+0129 11990.931 GB

19:00 UT - Experiment auto-started OK (vk)

11:30 UT - sky camera is stuck (Lucia)

11:43 UT - Ke reports about internet interruptions between 8pm and 1am tonight. No problems until now. (Lucia)

Yg HOB+0123 10227.158 GB

19:00 UT - Experiment auto-started OK (vk)

05:45 UT - Monica.Rxmon restarted (Liza)

08:18 UT - tsys in channel 5 too high at ~900 (known problem) (Lucia)

11:36 UT - tsys in bands 5-8 quite high (200-500) (Lucia)

You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/a1504.txt · Last modified: 2015/02/09 12:33 by Lucia Plank